OpenStudy (anonymous):

do 1 and 0.99999999... really represent the same number?

OpenStudy (anonymous):

yes.

OpenStudy (a_clan):

Not really But are close enough for calculation purpose.

OpenStudy (anonymous):

depends on what level of math you're at???

OpenStudy (anonymous):

yes. here's the proof: 0.999... = 0.9 + 0.09 + 0.009 + ... this is a geometric series with first term 0.9 and common ratio 0.1, which converges to 0.9/(1 - 0.1) = 1

OpenStudy (anonymous):

which are the levels?

OpenStudy (anonymous):

@santistebanc: well James just gave the higher and ultimately correct level. but try explaining that to a little kid.

OpenStudy (anonymous):

if A = 0.99999... then A = 0.9 + 0.09 + 0.009 + 0.0009 + ... then 0.1A = 0.09 + 0.009 + 0.0009 + ... thus A - 0.1A = (0.9 + 0.09 + 0.009 + 0.0009 + ...) - (0.09 + 0.009 + 0.0009 + ...) thus 0.9A = 0.9 thus A = 1

OpenStudy (anonymous):

I'm a little kid but thanks to everyone, now I get it.

OpenStudy (anonymous):

another way to prove it is like this x = .9999999999... 10x = 9.99999999... (Multiply both sides by 10) (Now subtract x from each side) 10x-x = 9x 9.9999999... - .99999999... = 9 9x = 9 (Divide by 9) x = 1

OpenStudy (anonymous):

0.33333.....(= 1/3) * 3 = 0.9999...(1).

OpenStudy (anonymous):

Gosh, thought we should have at least 100 posts by now:-)